Calculate the concentration of an EDTA solution of which 24.22 mL were needed to titrate the Ca2+ ions present in a 50.00 mL solution containing 246.7 mg CaCO3. Express your answer in terms of (a) molar concentration of EDTA and (b) Ca titer (mg Ca per mL of EDTA)
Answer #1

In the estimation of calcium ions by EDTA , Ca2+ ion solution is pepared by dissolving CaCO3 in HCl.

CaCO3 + 2 HCl → CaCl2 + CO2 + H2O

Note that EDTA is a hexadentate ligand , it forms an octahedral complex with calcium ions. The reaction of EDTA ( Ethylene diammine tetra acetate ) with Ca2+ takes place as follows :

(EDTA)4- + Ca2+   →Ca2+( EDTA) 4-  

Ca-EDTA complex is very strong complex.

Now in the given problem :

Ca2+ solution is prepared by dissolving 246.7 mg   ( ie 0.2467 gm) CaCO3 in 50 ml ( ie 50x 10-3 litre).

Hence number of moles of CaCO3 =   weight / molar mass = 0. 2467 / 100 = 2.467 x 10-3 moles

Molarity of Ca2+ ions = number of moles / volume of solution in litre

                                   = 2.467 x 10-3 / ( 50x10-3)      = 2.467 / 50 = 0.04934 M

Volume of EDTA = 24.22 ml

                              EDTA        -----------------------------          Ca2+ ions

                               M1V1                        =                                         M2V2

                               M1 x 24.22   =   0.04934 x 50

                                  M1 = 0.04934 x 50 / 24.22 = 0.1018 M

Molar Concentration of EDTA = 0.1018 M

(b) Ca titer (mg Ca per ml of EDTA)

In the lab total 24.22 ml EDTA was used to titrate 246.7mg Ca

Hence 1 ml EDTA was used to titrate = 246.7/24.22 = 10.185 mg Ca

Hence Ca titer is = 10.185 mg Ca per ml of EDTA